[figure: 9.7 cm]

THE MAGNETIC PROPERTIES OF (Y_{1−x}Zr_{x})Co_{4}B COMPOUNDS

R. Tetean, V. Pop, E. Burzo, I.Gr. Deac Babes-Bolyai University, Faculty of Physics, RO 3400 Cluj-Napoca, Romania

The (Y_{1−x}Zr_{x})Co_{4}B compounds with x ≤ 0.8 crystallize in CeCo_{4}B -type structure having P 6/mmm space group in the whole studied composition range. The lattice parameters decrease when yttrium atoms are substituted by zirconium ones. The magnetic measurements were performed in the temperature range 5 - 900 K and external magnetic fields up to 9 T. The compounds are ferromagneticaly ordered. The spontaneous magnetizations at 5 K decrease and the Curie temperatures increase when yttrium is gradually substituted by zirconium. Above the Curie points, the thermal variations of magnetic susceptibilities obey a Curie-Weiss type behaviour. The mean effective magnetic moments per cobalt atom were determined. The magnetic behaviour was correlated with the structure and composition of the compounds.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : R.Tetean

Address for correspondence : Babes-Bolyai University, Faculty of Physics, RO-3400 Cluj-Napoca, Romania Tel: ++4064194315 ext:196, Fax: ++4064191906

Email address : rote@phys.ubbcluj.ro